You throw a baseball directly upward at time=0 at an initial speed of 14.5 m/s. what is the maximum height the ball reaches above where it leaves your hand?
1
Expert's answer
2012-10-05T07:56:17-0400
At the highest point:
V = 0, V = V0 - g * t, so t = V0 / g
The maximum
height the ball reaches is:
h = V0 * t - g * t^2 / 2
h = V0 * V0 / g - g *
(V0 / g)^2 / 2 = V0^2 / (2 * g)
h = 14.5^2 / (2 * 9.81) = 10.7 m